I have no experiences using a logarithm so I need help trying to figure out how this formula works. Any assistance is appreciated.
Earthquakes The Richter scale is used to determine
the intensity of earthquakes, which corresponds
to the amount of energy released. If an earthquake has
an intensity of x, its magnitude, as computed by the
Richter scale, is given by R(x)= log x/i0 , where I is
the intensity of a small, measurable earthquake.
(a) On July 26, 1963, an earthquake in Yugoslavia
had a magnitude of 6.0 on the Richter scale, and
on August 19, 1977, an earthquake in Indonesia
measured 8.0. Find the intensity x for each of
these earthquakes if I0 = 1
(b) How many times more intense was the Indonesian
earthquake than the Yugoslavian earthquake?

Answer by stanbon(75887)

If an earthquake has
an intensity of x, its magnitude, as computed by the
Richter scale, is given by R(x)= log x/io , where I is
the intensity of a small, measurable earthquake.
-------
(a) On July 26, 1963, an earthquake in Yugoslavia
had a magnitude of 6.0 on the Richter scale, and
on August 19, 1977, an earthquake in Indonesia
measured 8.0. Find the intensity x for each of
these earthquakes if Io = 1
Yugo Quake: 6 = log(x/1)
6 = log(x)
x = 10^6
x = 1,000,000
========================
Indo Quake: 8 = log(x/1)
8 = log(x)
x = 10*8
x = 100,000,000
=========================
(b) How many times more intense was the Indonesian
earthquake than the Yugoslavian earthquake?
Ans: A factor of 100
---------------------------
Cheers,
Stan H.